OpenStudy (anonymous):

Solve (x/3) + 2 > 8.

OpenStudy (anonymous):

\[\frac{x}{3}+2>8\] right?

OpenStudy (anonymous):

add 2, multiply by 3

OpenStudy (mayankdevnani):

\[\large \bf \frac{x}{3}+2>8\] \[\large \bf \frac{6+x}{3}>8\]

OpenStudy (mayankdevnani):

and solve for x ?

OpenStudy (anonymous):

x=6?

OpenStudy (anonymous):

THE CONFUSION

OpenStudy (mayankdevnani):

\[\large \bf 6+x>8 \times 3\] \[\large \bf 6+x>24\]

OpenStudy (mayankdevnani):

\[\large \bf x>24-6\] \[\large \bf x>18\]

OpenStudy (texaschic101):

(x/3) + 2 > 8 --- multiply everything by 3 x + 6 > 24 -- subtract 6 from both sides x > 24 - 6 x > 18

OpenStudy (mayankdevnani):

understood ?? @lolcat

OpenStudy (anonymous):

im still really confused on how you do these

OpenStudy (texaschic101):

when you have an inequality/equation with fractions, multiply the entire equation by the common denominator and it will get rid of the fractions...making it easier to solve

OpenStudy (mayankdevnani):

1) first,solve x/3 +2